y''-4y'+3y=9x3+4 y(0)=6 y'(0)=8


1
Expert's answer
2021-08-11T04:47:52-0400

Given Differential equation y'' - 4y'+3y = 9x3 + 4 with y(0) = 6, y'(0) = 8

Solution:

Auxiliary equation corresponding to this given Differential equation is r2- 4r+3=0

"\implies" (r-3)(r-1) =0

"\implies" r=3 and r=1

therefore, yh = c1e3x + c2ex

Now to find particular integral (particular solution) = yp

Let yp = Ax3+Bx2+Cx+D

"\implies" yp' = 3Ax2+2Bx+C

"\implies" yp'' = 6Ax+2B


plugging this into given differential equation, yp'' - 4yp'+3yp = 9x3 + 4

"\implies" (6Ax+2B) - 4( 3Ax2+2Bx+C) + 3(Ax3+Bx2+Cx+D) = 9x3 + 4

"\implies" 3Ax3- 12Ax2+3Bx2+6Ax-8Bx+3Cx+2B-4C+3D = 9x3 + 4

comparing both sides, we get A = 3, B = 12,C = 26, D = 28

hence yp = 3x3 + 12x2 + 26x + 28


the general solution of the given differential equation is y = yh + yp

y = c1e3x + c2ex + 3x3 + 12x2 + 26x + 28

y' = 3c1e3x + c2ex + 9x2 + 24x + 26

Also given y(0) = 6, y'(0) = 8, plugging these conditions we get

6 = c1 + c2 + 28 "\implies" c1 + c2 = -22

and 8 = 3c1 + c2 + 26 "\implies" 3c1 + c2 = -18

By solving we get, c1 = 2 & c2 = -24

The solution of the differential equation is y = 2e3x - 24ex + 3x3 + 12x2 + 26x + 28